# thanks: The first two authors contributed equally

### v1

@@ VI Conclusion
We have presented, as a proof of principle, a toy model to detect the horizon supertranslation memory. The set up involves an IMBBH emitting GWs detectable by LISA, in the vicinity of an SMBH suddenly endowed with the memory, which acts as a proxy for a detector. The sudden change in the spacetime metric surrounding the SMBH produces a sharp change in the frequency and frequency evolution of the GWs, which can be exploited to identify the occurrence of this memory. We refer the reader to Goncharov et al. (2024) for other possible mechanisms that could enable the detection of BMS-symmetries through GW memory in next generation (XG) detectors.
@@
  • The change in the proper length of the IMBBHs is measured using a near-horizon metric. Thus, ideally, ST should be recovered near the horizon. On the other hand, realistically, the binary system should at least be outside the ISCO of the SMBH. The ISCO for the ST Schwarzschild black hole is not in general close to its horizon. For a bald Schwarzschild black hole, it is at $3$ times its Schwarzschild radius, and this distance may not be ideal for the detection scenario considered in this work. Of course, the ISCO lies very close to the horizon of a rapidly spinning Kerr black hole. Such rapidly spinning SMBHs are indeed expected in nature Reynolds (2015). Nevertheless, we have also considered the metric (II.1), proposed by the authors Hawking et al. (2017), which is valid for any finite radial distance from the center of the SMBH, and calculated the memory and the corresponding Log-Bayes-Factors. These values remain similar to what’s presented in Figure 3 at order of magnitude even if we place the detectors outside the ISCO of a supertranslated SMBH.
@@ Acknowledgments
We thank K. G. Arun for his feedback on the manuscript. S. A. B. acknowledges useful comments from Sourabh Magare and Avinash Tiwari. The research of S. B. is supported by DST-SERB through MATRICS grant MTR/2022/000170. S.J.K. acknowledges support from SERB grants SRG/2023/000419 and MTR/2023/000086.

### v2

@@ VII Conclusion
We have presented, as a proof of principle, two models to detect the horizon supertranslation memory. The first set up involves an IMBBH emitting GWs detectable by LISA, in the vicinity of a non-rotating SMBH suddenly endowed with the memory, which acts as a proxy for a detector. The sudden change in the spacetime metric surrounding the SMBH produces a sharp change in the frequency and frequency evolution of the GWs, which can be exploited to identify the occurrence of this memory. We further considered another case where the SMBH is an extreme Kerr black hole and determined the frequency change of the GWs. We refer the reader to Goncharov et al. (2024) for other possible mechanisms that could enable the detection of BMS-symmetries through GW memory in next generation (XG) detectors.
@@ Acknowledgments
We thank K. G. Arun and Swastik Bhattacharya for their feedback on the manuscript. S. A. B. acknowledges useful comments from Sourabh Magare and Avinash Tiwari. S.B. acknowledges warm hospitality at IUCAA, Pune during which the work was initiated. S.B. thanks the participants of the International Conference of Gravity & Cosmology (ICGC, 2023) held at IIT Guwahati for their inputs on an initial version of this work. S.A.B is grateful to the LISA Consortium participants for useful comments during the presentation of this work in the nascent stage as an invited talk in LISA Community Call and to the LISA Social Media Team (LSMT+) for advertising this work on LISA social media plateforms. The research of S. B. is supported by DST-SERB through MATRICS grant MTR/2022/000170. S.J.K. acknowledges support from SERB grants SRG/2023/000419 and MTR/2023/000086.